OpenStudy (anonymous):

A 50.0ml- volume of 0.15M HBr is titrated with 0.25M KOH . Calculate the pH after the addition of 13.0ml of KOH.

OpenStudy (anonymous):

meq of HBr=50*0.15=7.5 meq of KOH=0.25*13=3.25 meq of HBr left after titration=4.25 ph=-log (4.25*10^-3)=3-log4.25=3-log 4 (Approx)=3-0.602=2.398.